OpenStudy (anonymous):

Solve |2y - 2| = y There should be a solution set. Solve |2y - 2| = y There should be a solution set. @Mathematics

OpenStudy (anonymous):

positive case 2y -2 =y y=2 negative case -2y + 2 = y 3y=-2 y = -2/3 ss = 2, -2/3

myininaya (myininaya):

y can't be negative |something| can't have a negative output

myininaya (myininaya):

just y=2 is solution

OpenStudy (anonymous):

oops ... forgot to check

OpenStudy (anonymous):

on no myininaya!

OpenStudy (anonymous):

y cannot be negative, but \[2y-2\] can be negative in which case \[|2y-2|=2-2y\]

OpenStudy (anonymous):

so you really do get both solutions \[y=2,y=\frac{2}{3}\]

OpenStudy (anonymous):

and you can check them both to see that they work!

myininaya (myininaya):

i was just checking his solutions i did not actually solve it

OpenStudy (anonymous):

next you will be saying it is a typo!

myininaya (myininaya):

no for real i didn't solve it lol

OpenStudy (anonymous):

that is what i say anyway...

myininaya (myininaya):

you are so evil

OpenStudy (anonymous):

thank you

myininaya (myininaya):

lol your welcome

OpenStudy (anonymous):

now that i look carefully the initial answer is wrong in any case.